Number 20 - C please. The answer is 19.04 but can u explain how
Queston 20 of 30 625 points Tim Hortons paid its line workers $10 per hour last year when the Consqgmer Instructions: Round your answers to two decimal places. Price Index iwas 100. Suppose over the past year, deflation occurred and the aggregate price level fell to 84 a. Tim Hortons must pay its workers s b. Tim Hortons must pay its workers c.If Tim Hortons keeps the wage foxed at $10 per hour, in real terms, ts workers get a this year in order to keep the real wage fixed at $10 this year if it wants to increase the real wage by 8 percent increase in wages References eBook &Resources Learning Objactive: 08-04 Calculate the inflation rate by three methods and adjust nominal variables into real variables using a price index Numeric Response Section Using Price Indexes TO
Answer #1

Answer:

A) Real wage = Nominal wage*100/CPI

Given Real Wage =$10

CPI =84

Nominal Wage= (Real wage * CPI)/100=(10*84)/100=$8.4

So Horton must pay $8.4

B)

Increase in Real wage=8%

The new real wage =(1+8%) *10=$10.8

So nominal wage =10.8*84/100=$9.072

So Horton must pay $9.072

C)

If nominal wage =$10

Then Real wage =Nominal wage *100/ CPI=10*100/84=$11.904

So increase in real wage=(11.904-10)/10=19.04%

So increase in real wage for worker is 19.04%